OpenStudy (anonymous):

If the length and width of rectangle A are 10% and 30% less,respectively, than the length and width of rectangle B, then the area of A is equal to what percent of the area of B

OpenStudy (anonymous):

it will be 63 % of area of B

OpenStudy (anonymous):

how did you do that

OpenStudy (anonymous):

let c and d be the length and width of B, area =c*d now length and width of A will be 0.9c and 0.7d

OpenStudy (anonymous):

area of b= 0.9c * 0.7d= 0.63 *c *d=0.63* area of B which is 63%

OpenStudy (anonymous):

How does this percentage change if the length and width are 10% and 30% greater respectively?

OpenStudy (anonymous):

greater means area of b = (1.1 *c) *(1.3*d) =1.43*c*d= 43% increase from the original value